# When higher-order interactions enhance synchronization: the case of the Kuramoto model

### v1

@@ Results
Next, we explore how limited resources allocated between pairwise and higher-order interactions influence synchronization. To model resource constraints, we assume a fixed budget of $J=40$ arbitrary units, representing energy, material, or financial cost. Each link incurs a cost of $1$ arbitrary unit, while triangles cost $c_{2}$ units, with $c_{2}\in\{1,3,5\}$. Resources are allocated proportionally: a fraction $b_{1}$ of the budget is dedicated to links, while the remaining fraction $1-b_{1}$ is allocated to triangles; both are then chosen randomly, ensuring the resulting hypergraph is $\mathcal{H}$-connected. This allocation models practical scenarios where different interactions types carry heterogeneous resource costs. Beyond structural costs, synchronization also depends on coupling strengths $K_{1}$, $K_{2}$—associated with links and triangles, respectively—which are typically determined by the system’s physics. Our goal is to identify the optimal allocation strategy across varying relative costs and a range of coupling strength configurations. Additionally, we investigated how the coherence of initial conditions—from partially aligned to fully disordered—impacts the effectiveness of different allocation strategies. To systematically explore these factors, we evaluated synchronization across a grid of $51\times 51=2,601$ experimental conditions, varying coupling strength distributions ($K_{1}$, $K_{2}$) and allocation ratios ($b_{1}$). Synchronization performance was quantified by the time-averaged Kuramoto order parameter $R$, averaged over $N_{\mathcal{H}}=500$ random realizations per experimental condition to capture variability in structure, frequencies, and initial states. This exploration was repeated $6$ times in total, varying both the relative cost of triangles ($c_{2}\in\{1,3,5\}$) and the coherence of initial conditions ($\vartheta_{0}^{\mathrm{max}}=\pi/2$, for more coherent starts and $\vartheta_{0}^{\mathrm{max}}=2\pi$ for fully incoherent starts). Figure 4 summarizes the results of this extensive simulation campaign.
@@
Within each panel, the horizontal axis represents the fraction of resources allocated to links, $b_{1}$, and the vertical axis represents their associated coupling strength, $K_{1}$, with the coupling strength for triangles set as $K_{2}=0.3-K_{1}$; the sum of the two strengths is kept constant to ensure meaningful comparison across conditions. The color of each pixel denotes the order parameter $R$, averaged over both the second half of simulations and the $500$ hypergraph realizations, with brighter colors indicating stronger phase coherence. The black line denote the optimal allocation proportion $b_{1}$ for each combination of coupling strengths $K_{1}$ and $K_{2}$.
@@
A key finding from this analysis is that the optimal allocation parameter $b_{1}$ is rarely exactly 0 or 1, i.e., dedicating all resources exclusively to either links or triangles is almost never optimal. Trivially, full allocation to links is optimal only when $K_{2}\approx 0$; similarly, allocating the entire budget to triangles is optimal only when $K_{1}\approx 0$. Across all panels, the optimal configuration consistently favors a combination of both interaction types. Notably, the optimal value of $b_{1}$ strongly depends on the relative coupling strengths $K_{1}$, $K_{2}$, with more resources typically allocated to the interaction type associated with the stronger coupling.
@@
These findings indicate that higher-order interactions complement pairwise couplings and can enhance synchronization. To further validate this, we repeated the analysis under different conditions: using a different coupling function for the $3$-body interactions Skardal and Arenas (2020), and inspecting the Kuramoto-Daido order parameter $R_{2}$ instead of the Kuramoto order parameter $R$. In all cases, the results consistently confirm that a combination of links and triangles is more effective for synchronization than relying on a single interaction type. These additional results are reported in the Supplementary Materials.
@@ Discussion
Recent work on nonlinear oscillatory systems has shown that higher-order interactions often lead to richer dynamics and generally make synchronization more difficult to achieve. Except for the case in which the system starts within the attraction basin of the synchronous state Zhang et al. (2024); von der Gracht et al. (2024), which higher-order interactions make more robust, the presence of higher-order interactions generally conspires against synchronization, while other dynamical patterns occur, such as cluster synchronization, slow-switching, and multi-stability Tanaka and Aoyagi (2011); Skardal and Arenas (2020); León et al. (2024). In this work, we performed an extensive numerical study of the higher-order Kuramoto model, taken as a paradigmatic example of oscillatory dynamics with higher-order interactions. The former point was confirmed, i.e., that higher-order interactions make the synchronous state more robust, while the latter was partially corroborated as strong higher-order interactions were found to hamper synchronization; however, we uncovered a counterintuitive phenomenon, namely, that weak higher-order interactions can enhance synchronization.
A thorough numerical analysis across diverse settings, spanning a wide range of structure topologies and parameter configurations, revealed that, on average, weak higher-order interactions strengthened the synchronized state. This finding was further supported by a cost analysis, which showed that a combination of both pairwise and higher-order (3-body) interactions yields the most effective topology for achieving synchronization. Because we have averaged over several hundreds of randomly generated higher-order topologies, our results are valid in general, though there might exist specific hypergraph structures that deviate from our findings. Due to the computational cost of our analyses, we focused on relatively small hypergraphs. We expect our conclusions to remain valid for larger systems with link and triangle densities equal to or greater than those considered here, while sparser configurations might produce different outcomes, and thus represent an interesting direction for future work. Another promising avenue would be to identify which higher-order topologies most effectively enhance synchronization, by extending the numerical network analysis in Coraggio and di Bernardo (2024) to the higher-order setting.
@@
In conclusion, the insights gained in this study offer guidance for the design and control of complex systems with higher-order interactions D’Souza et al. (2023). We believe these results open the way to further theoretical studies of this phenomenon and will contribute to a deeper understanding of the interplay between higher-order topology and nonlinear dynamics at a fundamental level.
@@ Methods
In the Kuramoto model, higher-order (non-pairwise) interactions naturally emerge from phase reduction Nakao (2016); Pietras and Daffertshofer (2019); Monga et al. (2019). In general, phase reduction theory establishes that it is possible to meaningfully associate a phase oscillator dynamics to a dynamical system in a periodic regime (i.e., whose asymptotic solution is a stable limit cycle) Kuramoto (1984); Winfree (1980); Kuramoto and Nakao (2019). For example, the Kuramoto model itself Kuramoto (1975) can be derived from the complex-valued Stuart-Landau equation Nakao (2014), which can be rewritten as a two-dimensional system admitting a stable limit cycle. In high-dimensional networked systems with weak coupling, a similar phase reduction yields a system of phase oscillators interacting via the underlying network topology Nakao (2016). The classical Kuramoto model is obtained through a first order approximation with respect to the coupling parameter Kuramoto (1975, 1984). However, when phase reduction is performed using approximations beyond the first order Ashwin and Rodrigues (2016); León and Pazó (2019), non-pairwise interactions emerge. As an example, consider three oscillators, $j$, $k$ and $l$, coupled so that both $j$ and $l$ are connected to $k$ but not to each other. While first-order interactions exist only between $j$ and $k$, and $l$ and $k$, a small interaction term appears between $j$ and $l$, when second-order phase reduction is performed555 Note that the term higher-order is used ambiguously in the literature, as it can refer both to non-pairwise interactions (many-body coupling) and to higher-order expansions in the coupling parameter. Here, we use higher-order to mean non-pairwise interactions exclusively. Bick et al. (2024).
@@
These non-pairwise terms are typically small corrections, of the order of the squared coupling strength, and thus do not drastically alter the dynamics, but improve the accuracy of the reduced description Sakaguchi (1990); León and Pazó (2019); nonetheless, these small higher-order terms can have significant effects near bifurcation points Bick et al. (2024) or in phenomena such as remote synchronization Okuda and Kuramoto (1991); Bergner et al. (2012).
@@
The Kuramoto model with higher-order interactions (i.e., non-pairwise interactions of the same order as pairwise ones) has been extensively studied. However, these higher-order Kuramoto models were typically not derived via phase reduction from high-dimensional oscillators coupled through higher-order interactions. Instead, they were constructed by adding higher-order coupling terms to the classical pairwise Kuramoto model Tanaka and Aoyagi (2011); Skardal and Arenas (2019, 2020); Lucas et al. (2020). Various forms of higher-order coupling have been proposed in these studies, raising the question of which forms arise naturally from phase reduction. León et al. León et al. (2024) rigorously derived a higher-order Kuramoto model starting from complex Stuart-Landau oscillators coupled via pairwise and higher-order interactions that respect certain symmetry properties. They found that, for the dynamics of the $j$-th oscillator, the naturally emerging $3$-body interaction takes the form [^6] More precisely, the derivation also produces a phase lag, as in the Kuramoto-Sakaguchi model Sakaguchi and Kuramoto (1986). However, since much of the literature focuses on behavior without phase lag, we focus on that setting here. $\sin(\vartheta_{k}+\vartheta_{l}-2\vartheta_{j})$. This result motivated our choice of the higher-order Kuramoto model, namely,
$\dot{\vartheta_{j}}=\omega_{j}+\frac{K_{1}}{\langle d^{(1)}\rangle}\sum_{k=1}^{N_{0}}A_{jk}^{(1)}\sin(\vartheta_{k}-\vartheta_{j})\\ +\frac{K_{2}}{2\langle d^{(2)}\rangle}\sum_{k=1}^{N_{0}}\sum_{l=1}^{N_{0}}A_{jkl}^{(2)}\sin(\vartheta_{k}+\vartheta_{l}-2\vartheta_{j}).$ (2)
@@
A further generalization of phase reduction for arbitrary hypergraph topologies and coupling functions revealed that also the harmonic considered by Skardal and Arenas [^7] The higher-order interaction terms studied by Skardal and Arenas allow to apply the Ott-Antonsen reduction Ott and Antonsen (2008), which would not be possible with the interaction term given by $\sin(\vartheta_{k}+\vartheta_{l}-2\vartheta_{j})$. in Skardal and Arenas (2020), namely $\sin(2\vartheta_{k}-\vartheta_{l}-\vartheta_{j})$, can emerge from phase reduction when the non-reduced system includes certain couplings León et al. (2025). For this reason, we performed our numerical study also on the following higher-order Kuramoto model:
$\dot{\vartheta_{j}}=\omega_{j}+\frac{K_{1}}{\langle d^{(1)}\rangle}\sum_{k=1}^{N_{0}}A_{jk}^{(1)}\sin(\vartheta_{k}-\vartheta_{j})\\ +\frac{K_{2}}{2\langle d^{(2)}\rangle}\sum_{k=1}^{N_{0}}\sum_{l=1}^{N_{0}}A_{jkl}^{(2)}\sin(2\vartheta_{k}-\vartheta_{l}-\vartheta_{j}).$ (3)
@@
Synchronization can be measured via the Kuramoto order parameter $R\coloneqq\left\lvert\frac{1}{N_{0}}\sum_{j=1}^{N_{0}}e^{i\vartheta_{j}}\right\rvert$, where $R\approx 1$ indicates synchronization, and lower values of $R$ correspond to increasing incoherence. To detect two-cluster synchronization, it is possible to use the Kuramoto-Daido order parameter $R_{2}\coloneqq\left\lvert\frac{1}{N_{0}}\sum_{j=1}^{N_{0}}e^{2i\vartheta_{j}}\right\rvert$; in that case, $R$ would be low, while $R_{2}$ approaches $1$. In our simulations, we also computed $R_{2}$ and found it generally lower than $R$, indicating that, on average, no significant clustering phenomena occur; these results are shown in the Supplementary Materials.
Lastly, we note that the choice of frequency distribution is arbitrary. We sampled frequencies from uniform distributions, although using unimodal distributions, such as Gaussian or Lorentzian ones, is also valid and commonly used, as it allows some analytical treatment Acebrón et al. (2005). Notably, it is known that a Kuramoto model with uniform frequency distribution exhibits a first-order phase transition, while a models with unimodal distributions Pazó (2005) (or unimodal distributions with a plateau in the center Basnarkov and Urumov (2007)) display the usual second-order transition. However, these differences arise in the thermodynamic limit with global (all-to-all) coupling, which differs significantly from the finite, sparse systems studied here. Indeed, we verified that using a Gaussian distribution centered at $0.15$ with variance $0.1$ yields qualitatively similar results to those obtained using the uniform distribution in $[0,0.3]$.
@@
$z_{j}^{(1)}\coloneqq\sum_{k=1}^{n}A_{jk}^{(1)}\sin(\vartheta_{k}-\vartheta_{j}),\quad z_{j}^{(2)}\coloneqq\sum_{k=1}^{N_{0}}\sum_{l=1}^{N_{0}}A_{jkl}^{(2)}\sin(\vartheta_{k}+\vartheta_{l}-2\vartheta_{j}).$
For $1$-hyperedges, we first note that $z^{(1)}=B^{(1)}\sin(-(B^{(1)})^{\mathsf{T}}\vartheta)$, where $B^{(1)}$ is the incidence matrix. Then, for $2$-hyperedges, we define the two auxiliary matrices $Y^{(2)}_{\mathrm{a}},Y^{(2)}_{\mathrm{s}}\in\mathbb{Z}^{N_{0}\times 3N_{2}}$; the former is used to assemble linear combinations of the phases and the latter to select specific combinations contributing to the dynamics of each oscillator. In $Y^{(2)}_{\mathrm{a}}$, given some $k\in\{1,\dots,N_{2}\}$, the $3$ columns from the (3k2)(3k-2)-th to the $3k$-th correspond to the $k$-th 22-hyperedge, say $\{l,m,p\}$; the first of these columns has $2$ in position $l$, 1-1 in positions $m$ and $p$, and zeros elsewhere; the second has $2$ in $m$, and 1-1 in $l$ and $p$; the third has $2$ in $p$, and 1-1 in $l$ and $m$. In $Y^{(2)}_{\mathrm{s}}$, the $3$ columns from the (3k2)(3k-2)-th to the $3k$-th also correspond to the $k$-th 22-hyperedge, say $\{l,m,p\}$; these three columns are different one-hot vectors with $2$ in positions $l$, $m$ and $p$, respectively, and zeros elsewhere. An example set of the auxiliary matrices $Y^{(2)}_{\mathrm{a}}$, $Y^{(2)}_{\mathrm{s}}$ for a hypergraph with $N_{0}=4$ vertices and hyperedges $\{1,2,3\}$, $\{1,3,4\}$ is
@@
Then, it is immediate to verify that $z^{(2)}=Y^{(2)}_{\mathrm{s}}\sin(-(Y^{(2)}_{\mathrm{b}})^{\mathsf{T}}\vartheta)$.
@@
$\dot{\vartheta}=\omega+\frac{K_{1}}{\langle d^{(1)}\rangle}B_{1}\sin\left(-(B_{1})^{\mathsf{T}}\vartheta\right)+\frac{K_{2}}{2\langle d^{(2)}\rangle}Y^{(2)}_{\mathrm{s}}\sin\left(-(Y^{(2)}_{\mathrm{a}})^{\mathsf{T}}\vartheta\right),$ (4)
where $\vartheta$ and $\omega$ are $N_{0}$-dimensional column vectors, collecting all $\vartheta_{j}$’s and $\omega_{j}$’s, respectively. Similar steps can be applied to rewrite (3) in vector form; details are provided in the Supplementary Materials.
@@
Here, we illustrate the setup of the numerical simulations performed in this work. We integrated the higher-order Kuramoto model using the forward Euler method with time step $0.05$ for $600$ time units. [^8] Initially, results were obtained with a Runge-Kutta IV explicit method and a smaller time step of $0.005$, but since outcomes matched those from the Euler methods we opted for the latter to reduce the high computational cost. Integrating for longer times, e.g., $2000$ time units, also did not affect the results. The order parameter RR—a measure of synchronization—shown in the Results section figures is the time average over the second half of the simulation; time averaging is necessary as, even when starting close to the synchronous state, heterogeneity in frequencies and interactions often prevent the solution from becoming stationary. Our results represent averages over random hypergraph topologies: for each condition, we sampled $300$ hypergraphs for Figure 3 and $500$ for Figure 4. All simulations were implemented in Matlab The MathWorks Inc. (2022).

### v2

@@ Results
Next, we explore how limited resources allocated between pairwise and higher-order interactions influence synchronization. To model resource constraints, we assume a fixed budget of $J=40$ arbitrary units, representing energy, material, or financial cost. Each link incurs a cost of $1$ arbitrary unit, while triangles cost $c_{2}$ units, with $c_{2}\in\{1,3,5\}$. Resources are allocated proportionally: a fraction $b_{1}$ of the budget is dedicated to links, while the remaining fraction $1-b_{1}$ is allocated to triangles: specifically, the numbers of links and triangles are $N_{1}=\lfloor Jb_{1}/c_{1}\rfloor$ and $N_{2}=\lfloor J(1-b_{1})/c_{2}\rfloor$, respectively; both links and triangles are then chosen randomly, ensuring the resulting hypergraph is $\mathcal{H}$-connected. This allocation models practical scenarios where different interactions types carry heterogeneous resource costs. Beyond structural costs, synchronization also depends on coupling strengths $K_{1}$, $K_{2}$—associated with links and triangles, respectively—which are typically determined by the system’s physics. Our goal is to identify the optimal allocation strategy across varying relative costs and a range of coupling strength configurations. Additionally, we investigated how the coherence of initial conditions—from partially aligned to fully disordered—impacts the effectiveness of different allocation strategies. To systematically explore these factors, we evaluated synchronization across a grid of $51\times 51=2{,}601$ experimental conditions, varying coupling strength distributions ($K_{1}$, $K_{2}$) and allocation ratios ($b_{1}$). Synchronization performance was quantified by the time-averaged Kuramoto order parameter $R$, averaged over $N_{\mathcal{H}}=500$ random realizations per experimental condition to capture variability in structure, frequencies, and initial states. This exploration was repeated $6$ times in total, varying both the relative cost of triangles ($c_{2}\in\{1,3,5\}$) and the coherence of initial conditions ($\theta_{0}^{\mathrm{max}}=\pi/2$, for more coherent starts and $\theta_{0}^{\mathrm{max}}=2\pi$ for fully incoherent starts). Figure 5 summarizes the results of this extensive simulation campaign.
@@
Within each panel, the horizontal axis represents the fraction of resources allocated to links, $b_{1}$, and the vertical axis represents their associated coupling strength, $K_{1}$, with the coupling strength for triangles set as $K_{2}=0.3-K_{1}$; the sum of the two strengths is kept constant to ensure meaningful comparison across conditions. The color of each pixel denotes the order parameter $R$, averaged over both the second half of simulations and the $500$ hypergraph realizations, with brighter colors indicating stronger phase coherence. The black line denote the optimal link allocation fraction $b_{1}$ for each combination of coupling strengths $K_{1}$ and $K_{2}$.
@@
A key finding from this analysis is that the optimal link allocation fraction $b_{1}$ is rarely exactly 0 or $1$, i.e., dedicating all resources exclusively to either links or triangles is almost never optimal. Trivially, full allocation to links is optimal only when $K_{2}\approx 0$; similarly, allocating the entire budget to triangles is optimal only when $K_{1}\approx 0$. Across all panels, the optimal configuration consistently favors a combination of both interaction types. Notably, the optimal value of $b_{1}$ strongly depends on the relative coupling strengths $K_{1}$, $K_{2}$, with more resources typically allocated to the interaction type associated with the stronger coupling.
@@
These findings indicate that higher-order interactions complement pairwise couplings and can enhance synchronization. To further validate this, we repeated the analysis under different conditions: using the (211)(2-1-1) interaction as a different coupling function for the $3$-body interactions Skardal and Arenas (2020), inspecting the Kuramoto-Daido order parameter $R_{2}$ instead of the Kuramoto order parameter $R$, drawing the oscillator frequencies randomly from a Gaussian distribution rather than a uniform one, and considering larger hyperegraphs of $N_{0}=100$ nodes. In all cases, the results consistently confirm that a combination of links and triangles is more effective for synchronization than relying on a single interaction type. These additional results are reported in the Supplementary Materials.
@@ Discussion
Recent work on nonlinear oscillatory systems has shown that higher-order interactions often lead to richer dynamics and generally hinder synchronization. Except when the system starts within the attraction basin of the synchronous state Zhang et al. (2024); von der Gracht et al. (2024), which higher-order interactions make more robust, the presence of higher-order interactions generally conspires against synchronization, while other dynamical patterns occur, such as cluster synchronization, slow-switching, and multi-stability Tanaka and Aoyagi (2011); Skardal and Arenas (2020); León et al. (2024).
In this work, we performed a numerical study of the higher-order Kuramoto model on random hypergraphs—taken as a paradigmatic example of oscillatory dynamics with higher-order interactions—spanning various parameter configurations. We confirmed that higher-order interactions make the synchronous state more robust. At the same time, while strong higher-order interactions were found to hamper synchronization, weak higher-order interactions can instead enhance it (i.e., increase the order parameter, on average), when added to pairwise ones. This conclusion was further supported by a cost analysis, which showed that a combination of both pairwise and higher-order (3-body) interactions yields the most effective topology for achieving synchronization. Since our results are obtained by averaging over hundreds of random hypergraphs, they hold on average; specific hypergraph structures may behave differently, as reflected in the order parameter distributions in Figure 4.
These findings are consistent with previous analytical results obtained in all-to-all settings, both for the (1,1,2)(1,1,-2) interaction Chen et al. (2025) and the (2,1,1)(2,-1,-1) one Skardal and Arenas (2020), where it was shown that synchronization is slightly enhanced for weak higher-order coupling, prior to the possible onset of bistability at larger coupling strengths, or of explosive phenomena in general Kuehn and Bick (2021). Related results have also been reported independently in recent preprints on hyperring topologies, showing that weak higher-order interactions Wang et al. (2025b), as well as combinations of pairwise and higher-order interactions Skardal et al. (2025), enhance the stability of twisted states while preserving the structure of the attraction basin. Taken together, our results and those reported in the literature complement the picture introduced by Zhang et al. Zhang et al. (2024), according to which higher-order interactions make attraction basins “deeper but smaller”, by showing that weak higher-order interactions first affect basin depth before influencing basin size.
Due to the computational cost of our analyses, we focused on random hypergraphs with up to $100$ nodes, which are, nonetheless, relevant for real-world applications (e.g. Menara et al. (2022); Simpson-Porco et al. (2013); Alderisio et al. (2017)). In light of previous analytical results for all-to-all configurations Skardal and Arenas (2020); Chen et al. (2025), it is plausible that the phenomena identified here persist in larger systems with link and triangle densities comparable to or higher than those considered in this work, whereas sparser configurations may require ad hoc investigations.
An in depth exploration of different higher-order topologies represents a natural direction for future work. Another promising avenue would be to identify which higher-order topologies most effectively enhance synchronization, by extending the numerical network analysis in Coraggio and di Bernardo (2024) to the higher-order setting. Analytical studies on general hypergraph topologies are also needed to develop a comprehensive theoretical framework, to identify possible counterexamples, and to assess whether different frequency distributions (and connectivity type, e.g., $1$-connected vs. $\mathcal{H}$-connected) affect these phenomena in the thermodynamic limit and in very-large-scale systems, as frequency distributions are known to markedly shape the dynamics Pazó (2005). While this remains an open problem, the self-consistency analysis Ichinomiya (2004); Restrepo et al. (2005) developed by Chen et al. for the all-to-all higher-order setting Chen et al. (2025) provides a promising starting point.
@@
In conclusion, the insights gained in this study offer guidance for the design and control of complex systems with higher-order interactions D’Souza et al. (2023). We expect these results to stimulate further theoretical investigations and to contribute to a deeper understanding of the interplay between higher-order topology and nonlinear dynamics at a fundamental level.
@@ Methods
In the Kuramoto model, higher-order (non-pairwise) interactions naturally emerge from phase reduction Nakao (2016); Pietras and Daffertshofer (2019); Monga et al. (2019). In general, phase reduction theory establishes that it is possible to meaningfully associate a phase oscillator dynamics to a dynamical system in a periodic regime (i.e., whose asymptotic solution is a stable limit cycle) Kuramoto (1984); Winfree (1980); Kuramoto and Nakao (2019). For example, the Kuramoto model itself Kuramoto (1975) can be derived from the complex-valued Stuart-Landau equation Nakao (2014), which can be rewritten as a two-dimensional system admitting a stable limit cycle. In high-dimensional networked systems with weak coupling, a similar phase reduction yields a system of phase oscillators interacting via the underlying network topology Nakao (2016).
The classical Kuramoto model is obtained through a first order approximation with respect to the coupling parameter Kuramoto (1975, 1984). However, when phase reduction is performed using approximations beyond the first order Ashwin and Rodrigues (2016); León and Pazó (2019); Gengel et al. (2020), non-pairwise interactions emerge. As an example, consider three oscillators, $j$, $k$ and $l$, coupled so that both $j$ and $l$ are connected to $k$ but not to each other. While first-order interactions exist only between $j$ and $k$, and $l$ and $k$, a small interaction term appears between $j$ and $l$, when second-order phase reduction is performed [^6] Note that the term higher-order is used ambiguously in the literature, as it can refer both to non-pairwise interactions (many-body coupling) and to higher-order expansions in the coupling parameter. Here, we use higher-order to mean non-pairwise interactions exclusively. Bick et al. (2024).
@@
These non-pairwise terms are typically small corrections, of the order of the squared coupling strength, and thus do not drastically alter the dynamics, but improve the accuracy of the reduced description Sakaguchi (1990); León and Pazó (2019); Mau et al. (2024);
nonetheless, these small higher-order terms can have significant effects near bifurcation points Bick et al. (2024) or in phenomena such as remote synchronization Okuda and Kuramoto (1991); Bergner et al. (2012).
@@
The Kuramoto model with higher-order interactions (i.e., non-pairwise interactions of the same order as pairwise ones) has been extensively studied. However, these higher-order Kuramoto models were typically not derived via phase reduction from high-dimensional oscillators coupled through higher-order interactions. Instead, they were constructed by adding higher-order coupling terms to the classical pairwise Kuramoto model Tanaka and Aoyagi (2011); Skardal and Arenas (2019, 2020); Lucas et al. (2020).
Various forms of higher-order coupling have been proposed in these studies, raising the question of which forms arise naturally from phase reduction. León et al. León et al. (2024) rigorously derived a higher-order Kuramoto model starting from complex Stuart-Landau oscillators coupled via pairwise and higher-order interactions that respect certain symmetry properties. They found that, for the dynamics of the $j$-th oscillator, the naturally emerging $3$-body interaction takes the form [^7] More precisely, the derivation also produces a phase lag, as in the Kuramoto-Sakaguchi model Sakaguchi and Kuramoto (1986). However, since much of the literature focuses on behavior without phase lag, we focus on that setting here. $\sin(\theta_{k}+\theta_{l}-2\theta_{j})$ which we refer to as the (1,1,2)(1,1,-2) interaction. This result motivated our choice of the higher-order Kuramoto model, namely,
$\dot{\theta}_{j}=\omega_{j}+\frac{K_{1}}{\langle d^{(1)}\rangle}\sum_{k=1}^{N_{0}}A_{jk}^{(1)}\sin(\theta_{k}-\theta_{j})\\ +\frac{K_{2}}{2\langle d^{(2)}\rangle}\sum_{k=1}^{N_{0}}\sum_{l=1}^{N_{0}}A_{jkl}^{(2)}\sin(\theta_{k}+\theta_{l}-2\theta_{j}).$
@@
A further generalization of phase reduction for arbitrary hypergraph topologies and coupling functions revealed that also the term $\sin(2\theta_{k}-\theta_{l}-\theta_{j})$—i.e., the (2,1,1)(2,-1,-1) interaction considered [^8] The (2,1,1)(2,-1,-1) interaction studied, e.g., by Skardal and Arenas Skardal and Arenas (2020), allows to apply the Ott-Antonsen reduction Ott and Antonsen (2008), which would not be possible with the (1,1,2)(1,1,-2) interaction. in Skardal and Arenas (2020)—can emerge from phase reduction when the non-reduced system includes certain couplings León et al. (2025). For this reason, we performed our numerical study also on the following higher-order Kuramoto model:
$\dot{\theta}_{j}=\omega_{j}+\frac{K_{1}}{\langle d^{(1)}\rangle}\sum_{k=1}^{N_{0}}A_{jk}^{(1)}\sin(\theta_{k}-\theta_{j})\\ +\frac{K_{2}}{2\langle d^{(2)}\rangle}\sum_{k=1}^{N_{0}}\sum_{l=1}^{N_{0}}A_{jkl}^{(2)}\sin(2\theta_{k}-\theta_{l}-\theta_{j}).$ (2)
@@
Synchronization can be measured via the Kuramoto order parameter $R\coloneqq\left\lvert\frac{1}{N_{0}}\sum_{j=1}^{N_{0}}e^{i\theta_{j}}\right\rvert$, where $R\approx 1$ indicates synchronization, and lower values of $R$ correspond to increasing incoherence. To detect two-cluster synchronization, it is possible to use the Kuramoto-Daido order parameter $R_{2}\coloneqq\left\lvert\frac{1}{N_{0}}\sum_{j=1}^{N_{0}}e^{2i\theta_{j}}\right\rvert$; in that case, $R$ would be low, while $R_{2}$ approaches $1$. In our simulations, we also computed $R_{2}$ and found it generally lower than $R$, indicating that, on average, no significant clustering phenomena occur; these results are shown in the Supplementary Materials.
Lastly, we note that the choice of frequency distribution is arbitrary. We sampled frequencies from uniform distributions, although using unimodal distributions, such as Gaussian or Lorentzian ones, is also valid and commonly used, as it allows some analytical treatment Acebrón et al. (2005). Notably, it is known that a Kuramoto model with unimodal distributions displays the usual second-order transition to synchronization, while a model with uniform frequency distribution Pazó (2005) (or unimodal distributions with a plateau in the center Basnarkov and Urumov (2007)) exhibits a first-order phase transition. However, these differences arise in the thermodynamic limit with global (all-to-all) coupling, which differs significantly from the finite, sparse systems studied here. Indeed, in the Supplementary Materials we report simulations obtained using a Gaussian distribution centered at $0.15$ with standard deviation $0.1$, observing qualitatively similar results to those obtained using the uniform distribution in $[0,0.3]$.
@@
$z_{j}^{(1)}\coloneqq\sum_{k=1}^{n}A_{jk}^{(1)}\sin(\theta_{k}-\theta_{j}),\quad z_{j}^{(2)}\coloneqq\sum_{k=1}^{N_{0}}\sum_{l=1}^{N_{0}}A_{jkl}^{(2)}\sin(\theta_{k}+\theta_{l}-2\theta_{j}).$
For $1$-hyperedges, we first note that $z^{(1)}=B^{(1)}\sin(-(B^{(1)})^{\mathsf{T}}\theta)$, where $B^{(1)}$ is the incidence matrix. Then, for $2$-hyperedges, we define the two auxiliary matrices
$Y^{(2)}_{\mathrm{a}},Y^{(2)}_{\mathrm{s}}\in\mathbb{Z}^{N_{0}\times 3N_{2}}$; the former is used to assemble linear combinations of the phases and the latter to select specific combinations contributing to the dynamics of each oscillator. In $Y^{(2)}_{\mathrm{a}}$, given some $k\in\{1,\dots,N_{2}\}$, the $3$ columns from the (3k2)(3k-2)-th to the $3k$-th correspond to the $k$-th 22-hyperedge, say $\{l,m,p\}$; the first of these columns has $2$ in position $l$, 1-1 in positions $m$ and $p$, and zeros elsewhere; the second has $2$ in $m$, and 1-1 in $l$ and $p$; the third has $2$ in $p$, and 1-1 in $l$ and $m$. In $Y^{(2)}_{\mathrm{s}}$, the $3$ columns from the (3k2)(3k-2)-th to the $3k$-th also correspond to the $k$-th 22-hyperedge, say $\{l,m,p\}$; these three columns are different one-hot vectors with $2$ in positions $l$, $m$ and $p$, respectively, and zeros elsewhere. An example set of the auxiliary matrices $Y^{(2)}_{\mathrm{a}}$, $Y^{(2)}_{\mathrm{s}}$ for a hypergraph with $N_{0}=4$ vertices and hyperedges $\{1,2,3\}$, $\{1,3,4\}$ is
@@
Then, it is immediate to verify that $z^{(2)}=Y^{(2)}_{\mathrm{s}}\sin(-(Y^{(2)}_{\mathrm{b}})^{\mathsf{T}}\theta)$.
@@
$\dot{\theta}=\omega+\frac{K_{1}}{\langle d^{(1)}\rangle}B_{1}\sin\left(-(B_{1})^{\mathsf{T}}\theta\right)+\frac{K_{2}}{2\langle d^{(2)}\rangle}Y^{(2)}_{\mathrm{s}}\sin\left(-(Y^{(2)}_{\mathrm{a}})^{\mathsf{T}}\theta\right),$ (3)
where $\theta$ and $\omega$ are $N_{0}$-dimensional column vectors, collecting all $\theta_{j}$’s and $\omega_{j}$’s, respectively. Similar steps can be applied to rewrite (2) in vector form; details are provided in the Supplementary Materials.
@@
Here, we illustrate the setup of the numerical simulations performed in this work. We integrated the higher-order Kuramoto model using the forward Euler method with time step $0.05$ for $600$ time units. [^9] Initially, results were obtained with a Runge-Kutta IV explicit method and a smaller time step of $0.005$, but since outcomes matched those from the Euler methods we opted for the latter to reduce the high computational cost. Integrating for longer times, e.g., $2000$ time units, also did not affect the results. The order parameter RR—a measure of synchronization—shown in the Results section figures is the time average over the second half of the simulation; time averaging is necessary as, even when starting close to the synchronous state, heterogeneity in frequencies and interactions often prevent the solution from becoming stationary.
Our results represent averages over random hypergraph topologies: for each condition, we sampled $300$ hypergraphs for Figure 3 and $500$ for Figure 5. All simulations were implemented in Matlab The MathWorks Inc. (2024).